OpenStudy (anonymous):

Aki's Bicycle Designs has determined that when x hundred bicycles are built, the average cost per bicycle is given by C(x)=0.5x^2 -1.3x+9.972, where C(x) is in hundreds of dollars. How many bicycles should the shop build to minimize the average cost per bicycle?

OpenStudy (anonymous):

The shop should build ___ bicycles.

OpenStudy (anonymous):

do you remember formula for vertex?

OpenStudy (anonymous):

i wrote it maybe 15 times. use it here

OpenStudy (anonymous):

take the derivative and you can find the minimum by setting it equal to zero

OpenStudy (anonymous):

here we go...

OpenStudy (dumbcow):

C(x) represents a parabola where vertex is minimum Find vertex of C(x) using: x = -b/2a

OpenStudy (dumbcow):

haha satellite is that what that thing was about...lol

OpenStudy (anonymous):

yes and you will notice that first suggestion was... take the derivative!

OpenStudy (anonymous):

in this case it is a little odd because \[-\frac{b}{2a}=\frac{1.3}{1}=1.3\] oh i see in hundreds. should build 130

OpenStudy (dumbcow):

yes that would be easiest but im learning not to assume anyone knows calculus right off

OpenStudy (anonymous):

no in fact it would not be easiest!

OpenStudy (anonymous):

hilarious

OpenStudy (anonymous):

minus b over two a is easiest. why take the derivative, set equal zero and solve when you are going to get minus b over 2 a in any case?

OpenStudy (anonymous):

ikr? take the derivative. like i need to compute a limit to do this

OpenStudy (dumbcow):

lol either way..to me its easier